OpenStudy (anonymous):

If the electric potential in a region is given by V(x)=6/x2, the x component of the electric field in that region is? ... I'm not sure if it's −12x^(−3) or 12x^(-3). The given formula is Ex=-dV/dx but an example in the book doesn't multiply with the minus. (all help is really well appreciated)

OpenStudy (irishboy123):

convention has [we are in 1d here, i think]: \(\mathbf{E(x)} = - \dfrac{d}{dx} V(x)\) so if \( \large V(x)=\frac{6}{x^2}\), then \(\large -\frac{dV}{dx}\) might just come out as you expect it might.